[Brahms-l] Shift report 20050519 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu May 19 2005 - 08:01:09 EDT
Shifter leader was rk
Shifters were 


1. Very peaceful night with two long stores, one cont'd from previous shifted ended at 0440, the 2nd one still running started at 0540.
2. Runs 14755-14759 and 14762-14765 - data runs (FS@2.3deg@3450B, MRS@90deg@12B), 14760,14761 - pedestal runs.
3. None observed.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu May 19 08:01:48 2005

This archive was generated by hypermail 2.1.8 : Thu May 19 2005 - 08:01:55 EDT